# Functional bottlenecks can emerge from non-epistatic underlying traits

### v1

@@ Acknowledgements.
Figure S2: Linearization of the phenotypes according to the expansion described in Eq. (S2). (a) Scatter plot of the phenotype values shown in the main text, $E$, and after the expansion, $E^{(1)}$. The Spearman correlation coefficient is about $0.7$. The red and blue stars correspond respectively to the red and blue wild types. (b) Topology of paths surviving at the connectivity thresholds $E^{B}_{C}$ and $E^{R}_{C}$. Here $E$ is the linearized additive phenotype.
From now on, we stick to the choice made in the main text, of normalizing the enrichment scores to the wild types. We note that even after the non-linear fitness function has been inverted by the fitting procedure we described, the effect of epistasis is still present in the data. This is due to the fact that the fitness cannot be simply expressed as a nonlinear function of an additive trait. An interesting test is then to try and remove this second source of epistasis (network epistasis) and check whether one can still obtain a topology similar to the one discussed in Refs. Poelwijk et al. (2019); Alqatari and Nagel (2024) and in the main text. We do so by defining an additive phenotype $E^{(1)}_{f}$ obtained expanding the phenotypes $E_{f}$ around their respective wild type $\textbf{a}_{f}$ and cutting the higher order terms, namely
@@
Fig. S2a shows a scatter plot of the approximated additive phenotype $E^{(1)}(\textbf{a})$ against $E(\textbf{a})$. In the figure we highlight the wild types (blue and red stars) along with the variants that are at a small Hamming distance from the blue ($d_{B}$) and red ($d_{R}$) wild types. One notices that $E$ and $E^{(1)}$ are not simply linearly related, even if they have a relatively high Spearman correlation coefficient of $0.7$. This suggests that the first order expansion is not able to fully capture all mutational effects and that network epistasis is indeed present in the data Poelwijk et al. (2019). We note that the variants at small Hamming distance from the wild types display a better correlation than the very distant ones, as expected.
Keeping in mind that the linearization is only an approximation, we analyzed the space of viable paths connecting the two wild types. Following the same approach described in the main text, we look for the largest value of $E_{\rm th}$ for which at least one path allowing for a functionality switch survives, and call this value $E_{C}$. However, because the two wild types in this case turn out to be quite asymmetric, with $E^{WT}_{B}\sim 0.49$ and $E^{WT}_{R}\sim 2.8$, we allowed for an asymmetric connectivity threshold Alqatari and Nagel (2024) for the paths joining the two variants (see also the discussion in Sec. S1.2). Instead of increasing $E_{\rm th}$ symmetrically from 0 and impose $|E|>E_{\rm th}$ for every path, we impose $E>E^{B}_{\rm th}$ and $E<-E^{R}_{\rm th}$ separately for the two phenotypes. The largest values for which a path survives are $E^{B}_{C}=0.21$ and $E^{R}_{C}=1.64$, and we obtain the topology shown in Fig. S2b. Note that in the end, $E^{B}_{C}/E^{WT}_{B}\sim 0.42$ and $E^{R}_{C}/E^{WT}_{R}\sim 0.59$, hence we always end up with a connectivity threshold at about half the wild type phenotype. We observe a significant reduction in functional variants, approximately at half distance between the two wild types, and the emergence of a bottleneck topology.
@@
In Ref. Buda et al. (2023) an extensive review of combinatorial experiments similar to that of Ref. Poelwijk et al. (2019) is presented. The authors analyzed the results from ten studies (see references in Buda et al. (2023)) of combinatorially complete fitness landscapes of seven different enzymes, totaling 1440 genotype-phenotype data-points. The experimental results for fitness are then translated into an underlying phenotype by attempting at removing non-linearity (see Ref. Buda et al. (2023) for details), and the authors provide the resulting distribution of SMEs that can be downloaded from the paper repository.
@@
Figure S4: Statistical properties of the space of paths upon varying the number of mutations $M$, at fixed values of $p$, for the Gaussian (a),(b), and the Pareto cutoff (c),(d) distributions. (a),(c) The average value of $E_{C}/E^{WT}_{\max}$. (b),(d) The fraction of single jumpers present at $E_{C}$.
Figure S5: Statistics of a few relevant quantities for the model with Gaussian distribution, for $L=100,500,1000$. (a) Distribution of the number of mutations. (b) Distribution of the value of the positive (blue) wildtype divided by $E_{T}$. (c) Distribution of the selected SMEs $\tilde{h}$. (d) Distribution of $E_{C}/E^{WT}_{\max}$.
S3 Additional results on the toy model
@@
In the main text we presented the statistics of $E_{C}/E^{WT}_{\max}$ for the calibrated models, i.e. for the parameter choices $p$ and $E_{T}$ that maximize both the connectivity threshold $E_{C}/E^{WT}_{\max}$ and the probability of observing bottleneck structures. These two quantities are shown in Fig. S4 as a function of the number of mutations $M$ separating the wild types, for different values of $p$ and $E_{T}$. As explained in the main text, the dotted line at $E_{C}=0.5E^{WT}_{\max}$ corresponds to the maximum value of connectivity threshold that can be found when observing a bottleneck with a single functionality switch. At large $p$, it can be seen that although the fraction of bottlenecks remains equal to one, $\langle E_{C}/E^{WT}_{\max}\rangle$ decreases with increasing $p$. Low values of $E_{C}/E^{WT}_{\max}$ correspond to a connectivity threshold that lies deep within the region of non-functionality, making the corresponding evolutionary trajectories unlikely to occur. On the other hand, for small $p$, corresponding to a random selection of mutational effects, the fraction of observed bottlenecks decreases significantly.
For the values of $p$ and $E_{T}$ chosen after calibration (black dots), we observe simular results to Ref. Alqatari and Nagel (2024). The probability of a bottleneck remains always equal to one, while the value of $E_{C}/E^{WT}_{\max}$ increases with $M$ and saturates to the expected maximum.
Figure S6: Distribution of the relevant quantities for the model with Pareto distributed $h$ values. Results are shown for different exponents $\alpha=2.2$, $3.0$, and $5.0$. (a) Distribution of $E^{WT}_{B}/E_{T}$. Inset: tails of the distributions in log-log scale. The dashed lines are obtained fitting with an exponent equal to $\alpha$. (b) Distribution of the value of $E_{C}/E^{WT}_{\max}$. (c) Fraction of single jumpers at $E_{C}$ as a function of the number of mutations $M$.
@@
All the results reported in the main text have been obtained by fixing the number of SMEs to $L=500$. This would correspond to a protein of $500$ amino acids or to an elastic network with $500$ edges. To check the consistency of our results we also perform simulations varying the length $L$ of the binary sequence. In particular we study $L=100$ and $L=1000$ as well. We focus on the same parameters used in the main text, namely a Gaussian with $\mu=0$, $\sigma=1.0$ and we determine $p$, $E_{T}$ fixing $\langle M\rangle\approx 8$. As expected, the required value of $E_{T}$ increases with $L$, because the maximum of the $h_{i}$ increases proportionally to $\sqrt{\log L}$ as dictated by extreme value statistics. The growth is slow enough, however, and it is compensated by the tuning of $E_{T}$, in such a way that the results discussed in the main text are not significantly affected by the choice of the system size. In Fig. S5, panels (a) and (b), we plot the distribution of the number of mutations and of the value of $E^{WT}_{B}/E_{T}$, respectively. We see that, with an appropriate choice of $E_{T}$ and $p$ for each $L$, the results are almost indistinguishable for the three values of $L$. In panels (c) we plot the distribution of the values of $h_{i}$ selected by the tuning procedure. We see that, as $L$ increases, the peak of strongly beneficial (or strongly deleterious) mutations shifts towards larger absolute values, sharpening the separation between greedy and random steps. Finally, in panel (d) we show the distribution of $E_{C}/E^{WT}_{\max}$, which is once again unaffected by the choice of $L$.
@@
Yet, it is interesting to check what would happen in our model if the heavy tail would be allowed to continue up to infinity (keeping in mind that the finite value of $L$ will induce a cutoff anyways). Hence, we show here some results obtained for various choices of the input distribution $P(h)$ with Pareto tails decaying according to an exponent $\alpha$ as in Eq.(5) of the main text, without cutoff. We analyze different values of $\alpha=2.2,3,5$, for which the variance of the distribution is finite. We fix the system size to $L=500$ as in the main text, to facilitate comparison. We then proceed as described in the main text and we characterize the statistics of $E^{WT}_{{f}}/E_{T}$, $E_{C}/E^{WT}_{\max}$, and the topologies that arise from this input distribution of SMEs. The results are shown in figure S6, for the optimal choice of $p$ and $E_{T}$.
@@
For this choice of the parameters we find $\langle E_{C}/E^{WT}_{\max}\rangle\simeq 0.5$ which is the largest possible value of $E_{C}$ maximizing the fraction of bottleneck topologies in the ensemble. We note that if there is only one greedy step, the value of the wild type $E_{B}^{WT}$ is given by the sum of $M-1$ values drawn from $P(h)$ and a value drawn from $P^{L}_{\max}(h_{\max})$, the probability distribution of the maximum value of the $L$ variables $h$, which is easily derived from extreme value statistics. If $P(h)$ has heavy tails, we expect the sum to be dominated by the maximum and, as a consequence, $P(E^{WT}_{B})\sim P^{L}_{\max}(h_{\max})$. The same reasoning works for the red wild type $E^{WT}_{R}$. In Fig. S6a we show the plot of $E^{WT}_{B}$ for different values of the exponent $\alpha$. In the inset of this panel one can see how the tails of such distribution decay following a power law with the same exponent $\alpha$, as expected.
@@
The distribution of $E_{C}/E^{WT}_{\max}$, shown in Fig. S6b, seems to be rather independent of the exponent $\alpha$, and is centered around $0.5$ as observed in the previous cases. What is different is the fraction of single jumpers at $E_{C}$, shown in Fig. S6c. At variance with the Gaussian and Pareto cutoff cases, here the fraction of bottleneck topologies is different from one even for the optimal values of $p$ and $E_{T}$. This difference is due to the fat tails distribution of the wild type phenotype $E^{WT}_{{f}}$. Since the distribution is very broad, it is possible to have a very unbalanced situation in which one of the two wild types is much larger than the other (in particular, when $E^{WT}_{\max}\gtrsim 2E^{WT}_{\min}$), which allows for the saturation of $E_{C}$ to $E^{WT}_{\min}$ and the absence of a bottleneck. As expected, this effect disappears for larger values of $\alpha$, when the tails are less pronounced, and absent in the Gaussian and Pareto cutoff distributions.

### v2

@@ Acknowledgements.
Figure S2: Linearization of the phenotypes according to the expansion described in Eq. (S2). (a) Scatter plot of the phenotype values shown in the main text, $E$, and after the expansion, $E^{(1)}$. The Spearman correlation coefficient is about $0.7$. The red and blue stars correspond respectively to the red and blue reference variants. (b) Topology of paths surviving at the connectivity thresholds $E^{B}_{C}$ and $E^{R}_{C}$. Here $E^{(1)}$ is the linearized additive phenotype.
From now on, we stick to the choice made in the main text, of normalizing the enrichment scores to the ones of the reference variants. We note that even after the non-linear fitness function has been inverted by the fitting procedure we described, the effect of epistasis is still present in the data. This is due to the fact that the fitness cannot be simply expressed as a nonlinear function of an additive trait. An interesting test is then to try and remove this second source of epistasis (network epistasis) and check whether one can still obtain a topology similar to the one discussed in Refs. Poelwijk et al. (2019); Alqatari and Nagel (2026) and in the main text. We do so by defining an additive phenotype $E^{(1)}_{f}$ obtained expanding the phenotypes $E_{f}$ around their respective reference variant $\textbf{a}_{f}$ and cutting the higher order terms, namely
@@
Fig. S2a shows a scatter plot of the approximated additive phenotype $E^{(1)}(\textbf{a})$ against $E(\textbf{a})$. In the figure we highlight the reference sequences (blue and red stars) along with the ones that are at a small Hamming distance from the blue ($d_{B}$) and red ($d_{R}$) variants. One notices that $E$ and $E^{(1)}$ are not simply linearly related, even if they have a relatively high Spearman correlation coefficient of $0.7$. This suggests that the first order expansion is not able to fully capture all mutational effects and that network epistasis is indeed present in the data Poelwijk et al. (2019). We note that the variants at small Hamming distance from the reference ones display a better correlation than the very distant ones, as expected.
Keeping in mind that the linearization is only an approximation, we analyzed the space of viable paths connecting the two reference variants. Following the same approach described in the main text, we look for the largest value of $E_{\rm th}$ for which at least one path allowing for a functionality switch survives, and call this value $E_{C}$. However, because the two reference sequences in this case turn out to be quite asymmetric, with $E^{\rm{ref}}_{B}\sim 0.49$ and $E^{\rm{ref}}_{R}\sim 2.8$, we allowed for an asymmetric connectivity threshold Alqatari and Nagel (2026) for the paths joining the two variants (see also the discussion in Sec. S1.2). Instead of increasing $E_{\rm th}$ symmetrically from 0 and impose $|E|>E_{\rm th}$ for every path, we impose $E>E^{B}_{\rm th}$ and $E<-E^{R}_{\rm th}$ separately for the two phenotypes. The largest values for which a path survives are $E^{B}_{C}=0.21$ and $E^{R}_{C}=1.64$, and we obtain the topology shown in Fig. S2b. Note that in the end, $E^{B}_{C}/E^{\rm{ref}}_{B}\sim 0.42$ and $E^{R}_{C}/E^{\rm{ref}}_{R}\sim 0.59$, hence we always end up with a connectivity threshold at about half the reference phenotype. We observe a significant reduction in functional variants, approximately at half distance between the two reference sequences, and the emergence of a bottleneck topology.
@@
In Ref. Buda et al. (2023) an extensive review of combinatorial experiments similar to that of Ref. Poelwijk et al. (2019) is presented. The authors analyzed the results from ten studies (see references in Buda et al. (2023)) of combinatorially complete fitness landscapes of seven different enzymes, totaling 1440 genotype-phenotype data-points. The experimental results for fitness are then translated into an underlying phenotype by attempting at removing non-linearity (see Ref. Buda et al. (2023) for details), and the authors provide the resulting distribution of SMEs that can be downloaded from the paper repository. It should be noted, however, that most of these landscapes connect a reference wild type to an evolved variant with increased fitness, rather than two variants with distinct functionality.
@@
Figure S4: Statistical properties of the space of paths upon varying the number of mutations $M$, at fixed values of $p$, for the Gaussian (a),(b), and the Pareto cutoff (c),(d) distributions. (a),(c) The average value of $E_{C}/E^{\rm{ref}}_{\max}$. (b),(d) The fraction of single jumpers present at $E_{C}$.
Figure S5: Statistics of a few relevant quantities for the model with Gaussian distribution, for $L=100,500,1000$. (a) Distribution of the number of mutations. (b) Distribution of the value of the positive (blue) reference phenotype divided by $E_{T}$. (c) Distribution of the selected SMEs $\tilde{h}$. (d) Distribution of $E_{C}/E^{\rm{ref}}_{\max}$.
S3 Additional results on the stylized model
@@
In the main text we presented the statistics of $E_{C}/E^{\rm{ref}}_{\max}$ for the calibrated models, i.e. for the parameter choices $p$ and $E_{T}$ that maximize both the connectivity threshold $E_{C}/E^{\rm{ref}}_{\max}$ and the probability of observing bottleneck structures. These two quantities are shown in Fig. S4 as a function of the number of mutations $M$ separating the reference variants, for different values of $p$ and $E_{T}$. As explained in the main text, the dotted line at $E_{C}=0.5E^{\rm{ref}}_{\max}$ corresponds to the maximum value of connectivity threshold that can be found when observing a bottleneck with a single functionality switch. At large $p$, it can be seen that although the fraction of bottlenecks remains equal to one, $\langle E_{C}/E^{\rm{ref}}_{\max}\rangle$ decreases with increasing $p$. Low values of $E_{C}/E^{\rm{ref}}_{\max}$ correspond to a connectivity threshold that lies deep within the region of non-functionality, making the corresponding evolutionary trajectories unlikely to occur. On the other hand, for small $p$, corresponding to a random selection of mutational effects, the fraction of observed bottlenecks decreases significantly.
For the values of $p$ and $E_{T}$ chosen after calibration (black dots), we observe simular results to Ref. Alqatari and Nagel (2026). The probability of a bottleneck remains always equal to one, while the value of $E_{C}/E^{\rm{ref}}_{\max}$ increases with $M$ and saturates to the expected maximum.
Figure S6: Distribution of the relevant quantities for the model with Pareto distributed $h$ values. Results are shown for different exponents $\alpha=2.2$, $3.0$, and $5.0$. (a) Distribution of $E^{\rm{ref}}_{B}/E_{T}$. Inset: tails of the distributions in log-log scale. The dashed lines are obtained fitting with an exponent equal to $\alpha$. (b) Distribution of the value of $E_{C}/E^{\rm{ref}}_{\max}$. (c) Fraction of single jumpers at $E_{C}$ as a function of the number of mutations $M$.
@@
All the results reported in the main text have been obtained by fixing the number of SMEs to $L=500$. This would correspond to a protein of $500$ amino acids or to an elastic network with $500$ edges. To check the consistency of our results we also perform simulations varying the length $L$ of the binary sequence. In particular we study $L=100$ and $L=1000$ as well. We focus on the same parameters used in the main text, namely a Gaussian with $\mu=0$, $\sigma=1.0$ and we determine $p$, $E_{T}$ fixing $\langle M\rangle\approx 8$. As expected, the required value of $E_{T}$ increases with $L$, because the maximum of the $h_{i}$ increases proportionally to $\sqrt{\log L}$ as dictated by extreme value statistics. The growth is slow enough, however, and it is compensated by the tuning of $E_{T}$, in such a way that the results discussed in the main text are not significantly affected by the choice of the system size. In Fig. S5, panels (a) and (b), we plot the distribution of the number of mutations and of the value of $E^{\rm{ref}}_{B}/E_{T}$, respectively. We see that, with an appropriate choice of $E_{T}$ and $p$ for each $L$, the results are almost indistinguishable for the three values of $L$. In panels (c) we plot the distribution of the values of $h_{i}$ selected by the tuning procedure. We see that, as $L$ increases, the peak of strongly beneficial (or strongly deleterious) mutations shifts towards larger absolute values, sharpening the separation between greedy and random steps. Finally, in panel (d) we show the distribution of $E_{C}/E^{\rm{ref}}_{\max}$, which is once again unaffected by the choice of $L$.
Figure S7: Distribution of $P(\tilde{h}/\tilde{h}_{\rm{max}})$. Comparison between calibrated (bold $\bm{p}$) and non calibrated models. (a) Gaussian input distribution $P(h)$. Calibrated model for $p=0.26$, $E_{T}=2.0$. (b) Pareto cutoff distribution $P(h)$: $\alpha=0.7$, $x_{m}=1$, $\epsilon=2$. Calibrated model for $p=0.25$, $E_{T}=1.1$. (c) Pareto distribution $P(h)$ with $\alpha=2.0$. Calibrated model for $p=0.3$, $E_{T}=8.0$.
@@
Yet, it is interesting to check what would happen in our model if the heavy tail would be allowed to continue up to infinity (keeping in mind that the finite value of $L$ will induce a cutoff anyways). Hence, we show here some results obtained for various choices of the input distribution $P(h)$ with Pareto tails decaying according to an exponent $\alpha$ as in Eq.(5) of the main text, without cutoff. We analyze different values of $\alpha=2.2,3,5$, for which the variance of the distribution is finite. We fix the system size to $L=500$ as in the main text, to facilitate comparison. We then proceed as described in the main text and we characterize the statistics of $E^{\rm{ref}}_{{f}}/E_{T}$, $E_{C}/E^{\rm{ref}}_{\max}$, and the topologies that arise from this input distribution of SMEs. The results are shown in Fig. S6, for the optimal choice of $p$ and $E_{T}$.
@@
For this choice of the parameters we find $\langle E_{C}/E^{\rm{ref}}_{\max}\rangle\simeq 0.5$, which is the largest possible value of $E_{C}$ maximizing the fraction of bottleneck topologies in the ensemble. We note that if there is only one greedy step, the value of the reference phenotype $E_{B}^{\rm{ref}}$ is given by the sum of M1M-1 values drawn from $P(h)$ and a value drawn from $P^{L}_{\max}(h_{\max})$, the probability distribution of the maximum value of the $L$ variables $h$, which is easily derived from extreme value statistics. If $P(h)$ has heavy tails, we expect the sum to be dominated by the maximum and, as a consequence, $P(E^{\rm{ref}}_{B})\sim P^{L}_{\max}(h_{\max})$. The same reasoning works for the red reference phenotype $E^{\rm{ref}}_{R}$. In Fig. S6a we show the plot of $E^{\rm{ref}}_{B}$ for different values of the exponent $\alpha$. In the inset of this panel one can see how the tails of such distribution decay following a power law with the same exponent $\alpha$, as expected.
@@
The distribution of $E_{C}/E^{\rm{ref}}_{\max}$, shown in Fig. S6b, seems to be rather independent of the exponent $\alpha$, and is centered around $0.5$ as observed in the previous cases. What is different is the fraction of single jumpers at $E_{C}$, shown in Fig. S6c. At variance with the Gaussian and Pareto cutoff cases, here the fraction of bottleneck topologies is different from one even for the optimal values of $p$ and $E_{T}$. This difference is due to the fat tails distribution of the reference phenotype $E^{\rm{ref}}_{{f}}$. Since the distribution is very broad, it is possible to have a very unbalanced situation in which one of the two reference phenotypes is much larger than the other (in particular, when $E^{\rm{ref}}_{\max}\gtrsim 2E^{\rm{ref}}_{\min}$), which allows for the saturation of $E_{C}$ to $E^{\rm{ref}}_{\min}$ and the absence of a bottleneck. As expected, this effect disappears for larger values of $\alpha$, when the tails are less pronounced, and absent in the Gaussian and Pareto cutoff distributions.
S3.4 Balance between neutral and strongly non-neutral mutations
A central finding of our work is that the emergence of functional bottlenecks is contingent upon a specific balance within the distribution of a posteriori ‘fixed’ mutational effects, P(h~)P($\tilde${h}). While the a priori distribution of available mutations, $P(h)$, can be relatively flexible, the realized mutational neighborhood of the reference genotypes must exhibit a high degree of heterogeneity to produce the desired topology. In Fig. S7, we illustrate this balance by comparing calibrated models against non-calibrated ones. The calibration procedure essentially ensures that the two reference genotypes are separated by an underlying trait distance that allows for a ‘neutral bulk’ of mutations punctuated by a sufficient frequency of large-effect mutations. As illustrated in Fig. S7, when the model is not properly calibrated, the resulting distribution P(h~)P($\tilde${h}) fails to achieve this critical composition in two distinct ways. If the fraction of large-effect mutations is too small, the fitness landscape remains overly smooth and accessible, preventing the formation of fitness valleys (at too small $p$). Conversely, if too many large-effect mutations dominate the distribution (at too large $p$), the landscape becomes fragmented, and the selection pressure required to form a ‘corridor’ between functional states must be too small. These results substantiate our claim that functional bottlenecks do not emerge by default from any nonlinear mapping. Rather, they require a specific balance between a majority of nearly neutral mutations and a minority of strongly non-neutral ones, which is provided by our procedure to construct the reference genotypes. This provides a clear example of why the construction is non-trivial: the bottleneck structure is a direct consequence of the heterogeneity in mutational effect sizes, and it vanishes when this heterogeneity is removed or incorrectly tuned.
Figure S8: Topology of paths surviving at the connectivity threshold for two instances drawn from uncalibrated models. $E$ and $E_{C}$ are scaled with respect to the largest of the two reference genotypes $E^{{\rm{{ref}}}}_{\rm{max}}$. (a) Gaussian input distribution $P(h)$, $p=0.1$, $E_{T}=0.08$. Note the typical behaviour: the two reference genotypes are unbalanced, $E^{{\rm{{ref}}}}_{B}/E_{T}=2.73$, $E^{{\rm{{ref}}}}_{R}/E_{T}=-9.25$, $E^{{\rm{{ref}}}}_{B}/E^{{\rm{{ref}}}}_{R}=3.39$ ; the connectivity threshold reaches its upper bound, which is the smallest between the two reference genotypes $E^{{\rm{{ref}}}}_{\rm{min}}$. (b) Gaussian input distribution $P(h)$, $p=0.8$, $E_{T}=8.2$. Here the reference genotypes are symmetric and close to $E_{T}$: $E^{{\rm{{ref}}}}_{B}/E_{T}=1.11$, $E^{{\rm{{ref}}}}_{R}/E_{T}=-1.06$, $E^{{\rm{{ref}}}}_{B}/E^{{\rm{{ref}}}}_{R}=1.04$; there is a single jumper mutation at half distance but the connectivity threshold is very small.
In Fig. S8 we show typical examples of paths that emerge when the model is not properly calibrated. As already stated before, if the fraction of large-effect mutations is too small, no single jumper mutation stands out and the functionality switch can occur at any step along the evolutionary path. This is illustrated in Fig. S8a: the fitness landscape is smooth and no bottleneck emerges during evolution. Conversely, if large-effect mutations dominate the distribution, as in Fig. S8b, the jump required to switch functionality is small: evolution has to accommodate substantial functionality loss with respect to the reference genotypes before being able to switch functionality, thus making these paths very unlikely to occur.